# Learning Transferable Friction Models and LuGre Identification via Physics Informed Neural Networks

### v1

@@ IV-D Parameter Estimation Performance
Choosing an appropriate friction model for each setup is an important decision. The complexity of the friction model, the identification process for parameters, the time it takes, and the computational cost are all essential decision factors in deciding which friction model is the best for a particular setup. Although the LuGre friction model accurately captures most aspects of the actual frictional behavior, such as stick-slip, it requires more work to identify the parameters that fit the experimental data. The dynamic nature of the LuGre friction model requires the internal state to be estimated via the differential equations that define the system and the LuGre friction model. Due to the high resolution required to capture the transition regions (stick-slip), variable step solvers like ode45 are required. This lends itself to high computational cost and high computation times, often resulting in simpler friction models being chosen.
We show that methods employing PINNs lead to quicker identification of similarly complex and sufficiently accurate friction models (Table III). $PE_{1,2}$ models in Fig. 2b estimate the internal LuGre state $z$ from the model inputs, which removes the need for simulating the LuGre dynamics, resulting in quicker parameter estimation. Although this approach might not replace a full parameter identification in safety critical systems where accuracy is paramount, it is a quicker way of estimating LuGre parameters. Hence, our approach lies between a more time consuming full parameter identification of the LuGre friction model and use of simpler friction models (Fig. 8), providing a fast approach toward bridging the sim-to-real gap.
Figure 8: Qualitative comparison of computation time and accuracy of different friction parameter ID approaches .
@@ V Conclusion and Future Work
We formulated a PINN-based friction estimation framework for creating transferable learned friction models. We showed the efficacy of these models both as in-simulation models and online friction estimators. We also presented the transferability of these models on a system that they were not trained on. We then discussed our framework’s computational speed and accuracy in parameter identification of the LuGre friction model and the appeal of this approach as a fast way of shrinking the sim-to-real gap. Future work includes incorporating this framework with existing control methods, trajectory optimization, and expanding the framework to accommodate online adaption to changing friction properties. In each such scenario, we anticipate a Pareto trade-off of the same general form (Fig. 8), with the exact trade-offs between computational speed and accuracy of motion being a key concern for future work.

### v2

@@ IV-D Parameter Estimation Performance
Choosing an appropriate friction model for each setup is an important decision. The complexity of the friction model, the identification process for parameters, the time it takes, and the computational cost are all essential decision factors in deciding which friction model is the best for a particular setup. Although the LuGre friction model accurately captures most aspects of the actual frictional behavior, such as stick-slip, it requires more work to identify the parameters that fit the experimental data.
We show that our methods employing PINNs enable faster identification of friction models of comparable complexity with sufficient accuracy. Table III compares the parameter estimation performance of established methods with our two models $PE_{1,2}$ shown in Fig. 2b. The performance of our models is on par with the other methods. A performance trade-off between mean squared error (MSE) of the estimated friction forces and computation time is illustrated in Fig. 8. While our approach may not replace full parameter identification in safety-critical systems where maximum accuracy is essential, it provides a faster alternative for estimating LuGre parameters. Although Nelder-Mead simplex method appears competitive in Fig. 8, its parameter estimation accuracy was worse (Table III).
@@ V Conclusion and Future Work
We formulated a PINN-based friction estimation framework for creating transferable learned friction models. We demonstrated the efficacy of these models both as in-simulation friction models and online friction estimators, showed their transferability to a dynamical system they were not trained on, and compared the framework’s computational speed and accuracy in LuGre parameter identification against established methods.
The present results are validated using synthetic data generated from the LuGre model. While this demonstrates that the framework can recover complex friction behavior from minimal state measurements, experimental validation on physical hardware is needed to fully assess its applicability to real-world systems. Further work includes extending the framework to multi-contact and higher degree-of-freedom systems, systematic comparison with other hybrid learning-based friction modeling approaches, and integration with existing control methods and trajectory optimization. Expanding the framework to accommodate online adaptation to changing friction properties is also planned. In each such scenario, we anticipate a Pareto trade-off of the same general form (Fig. 8), with the exact trade-offs between computational speed and accuracy of motion being a key concern.